OpenStudy (anonymous):

What is the sum of the rational expressions below? 3x+1-2x+7x/x+5 A. 10x+1/3x+5 B. 17x^2+16x+5/2x^2+10x C. 3x^2+16x+5/2x^2+10x D.17x^2+16x+5/2x^2+5

OpenStudy (danjs):

retype it, something is wrong in the original expression (x/x)=1

OpenStudy (anonymous):

What is the sum of the rational expressions below? 3x+1/2x+7x/x+5

OpenStudy (danjs):

\[3x+\frac{ 1 }{ 2 }x+\frac{ 7x }{ x }+5\] that?

OpenStudy (anonymous):